OpenStudy (anonymous):

cos2x + 2 cos x + 1 = 0 Find all solutions in the interval [0, 2π).

OpenStudy (anonymous):

it is clear that this is the same thing as \[(\cos(x)+1)^2=0\]?

OpenStudy (anonymous):

making \(\cos(x)+1=0\) i.e. \(\cos(x)=-1\)

OpenStudy (anonymous):

Consider plotting this and see x values where it = 0